# Bayesian Controlled FDR Variable Selection via Knockoffs

### v2

@@ 2.2 Constructing Bayesian knockoffs
We consider a linear regression setting, where we aim at simultaneously inferring the subset of active predictors while also learning the conditional independence structure of the covariates. Thus, we specify the distribution $p(Y\mid X,\widetilde{X})$ as a linear regression model of the type
$Y=X\beta+\widetilde{X}\widetilde{\beta}+\epsilon,\qquad\epsilon\sim N(0,\sigma^{2}),$ (7)
where $\beta$ and $\widetilde{\beta}$ are $p$-dimensional vectors of the covariates’ effects and the knockoffs’ effects on $Y$, respectively. Following the decomposition introduced in Equation (4), we can rewrite the model in Equation (7) as
$Y=X\left[\beta+(I_{p}-\bm{\Omega}\text{diag}(s))\widetilde{\beta}\right]+U\widetilde{\beta}+\epsilon,\qquad\epsilon\sim N(0,\sigma^{2}),$ (8)
where $\bm{\Omega}=\bm{\Sigma}^{-1}$ is the precision matrix. Under this specification, it is possible to obtain the target distribution $p(Y\mid X)$ in closed form as
@@
with the matrix ${\bm{A}}$ as defined in Equation (2). This construction respects the conditional independence property of the knockoff variables, as now $Y\mbox{$\perp\!\!\!\perp\,$}\widetilde{X}\mid X$. Moreover, since the decomposition of the knockoff variables in Equation (4) is obtained from the joint distribution in Equation (1), our formulation respects the pairwise exchangeability too. Therefore, this setting uses valid model-X knockoffs. Note that the construction of the Bayesian knockoff filter proposed by Gu and Yin, (2021) does not produce valid knockoffs, since adopting $p(Y\mid X,\widetilde{X})$ as the likelihood function, as they do, implies $Y\mbox{$\perp\!\!\!\perp\!\!\!\!\!\!\setminus\;$}\widetilde{X},X$ and therefore the conditional independence property is not satisfied.
@@ 2.2.1 Sparsity inducing prior for covariates’ graphical model
We regard the Gaussian distribution (5) on the covariates as a graphical model and impose sparsity in the associated graph structure. Under the assumption of joint Gaussian distribution, conditional independence implies zero values in the off-diagonal elements of the precision matrix $\bm{\Omega}=\bm{\Sigma}^{-1}$. When $\bm{\Omega}$ is sparse, also the matrix $\bm{A}$ in Equation (2) is sparse, facilitating the sampling of the knockoff variables under the pairwise exchangeability property. In addition, promoting sparsity in the precision matrix $\bm{\Omega}$ greatly improves computational efficiency and improves practical interpretability across various application domains. Here, we achieve this by imposing a continuous spike-and-slab prior (Wang,, 2015) that combines a mixture prior distribution on the off-diagonal elements $\omega_{jj^{\prime}}$ with an exponential distribution on the diagonal ones $\omega_{jj}$ ($j=1,\ldots,p$). Let $\mathcal{G}$ be the undirected graph representing the conditional independences among the covariates, and let $\bm{G}$ be the adjacency matrix with elements $g_{jj^{\prime}}\in\{0,1\}$, for $j,j^{\prime}=1,\ldots,p$ and $j\neq j^{\prime}$, with $1$ representing the presence of an edge between nodes $(j,j^{\prime})$. Assuming the parameters in $\bm{\Omega}$ independent a priori, the prior distribution can be written as
@@ 2.2.2 Priors for variable selection
For the regression coefficients, we adapt the spike-and-slab prior introduced in the supplementary material of Candès et al., (2018). Specifically, we modify their formulation by allowing each variable’s inclusion probability to depend on its number of neighbors in the estimated covariate graph, as detailed below. For simplicity of exposition, here we refer to the conditional model in (7), although similar arguments apply to the marginal model (9). We first specify the joint distribution
@@
with $j=1,\ldots,p$. Here, $h_{\beta}$ is a positive scaling hyperparameter and $\gamma=(\gamma_{1},\dots,\gamma_{p})^{T}\in\{0,1\}^{p}$ is a $p$-dimensional vector of binary indicators such that $\gamma_{j}=1$ if $X_{j}$ or its knockoff $\widetilde{X}_{j}$ are included in the model and, alternatively, $\gamma_{j}=0$ if neither the original feature nor the knockoff is included. By introducing two additional latent indicator vectors, $\delta=(\delta_{1},\dots,\delta_{p})$ and $\widetilde{\delta}=(\widetilde{\delta}_{1},\dots,\widetilde{\delta}_{p})$, such that $\gamma=\delta+\widetilde{\delta}$, the joint prior in Equation (12) can be reformulated as a mixture of independent spike-and-slab priors on $\beta_{j}$ and $\tilde{\beta}_{j}$ as
@@
Including both a covariate and its knockoff simultaneously in the model would introduce bias in the regression coefficient estimates. Therefore, we do not allow both indicators to assume value $1$ and choose a multinomial distribution of the type
$\left(\delta_{j},\widetilde{\delta}_{j}\right)\mid\gamma$$\sim\begin{cases}(0,0)\quad\text{ w.p. }\mathbb{P}[\gamma_{j}=0]\\ (0,1)\quad\text{ w.p. }\mathbb{P}[\gamma_{j}=1]/2\\ (1,0)\quad\text{ w.p. }\mathbb{P}[\gamma_{j}=1]/2.\end{cases}$
Furthermore, we let the probability of success depend on the graphical model on the covariates, similar to Peterson et al., (2016), via an Ising prior of the type
@@
We note that Candès et al., (2018) adopt prior in Equation (12) in a fixed design Bayesian regression with augmented covariates matrix $[\bm{X},\widetilde{\bm{X}}]$ where $\widetilde{\bm{X}}$ is sampled once from Equation (3). On the contrary, we consider the knockoffs as latent variables and update them in the posterior sampling along with the other parameters.
@@ 2.2.3 Markov Chain Monte Carlo for posterior sampling
$\begin{split}&p(\Theta\mid\bm{D})\propto{}p(Y\mid X,\Theta)\;p(X\mid\Theta)\;p(\Theta)\propto{}\\ \propto{}&p(Y\mid X,\beta,\widetilde{\beta},\bm{\Omega},\sigma^{2})p(X\mid\bm{\Omega})p(\beta,\widetilde{\beta}\mid\delta,\widetilde{\delta})p(\delta,\widetilde{\delta}\mid\gamma)p(\bm{\Omega}\mid\bm{G})p(\gamma\mid\bm{G})p(\bm{G})p(\sigma^{2}).\\ \end{split}$
This distribution is not analytically tractable, and we therefore resort to a Markov Chain Monte Carlo (MCMC) algorithm (Brooks et al.,, 2011). In particular, we implement an efficient Metropolis within Gibbs scheme that employs a data augmentation (DA) scheme based on the latent variable $U$ introduced in (4) and the formulation of the joint model detailed in Section 2.2. Specifically, as it is not possible to obtain tractable full conditional distributions using the marginal model $p(Y\mid X,\beta,\widetilde{\beta},\bm{\Omega},\sigma^{2})$, we employ a DA step by sampling from the distribution $p(Y\mid X,U,\beta,\widetilde{\beta},\bm{\Omega},\sigma^{2})$ defined in Equation (8) and then updating $U$ from the full conditional. We also incorporate the Add-Delete Metropolis-Hastings algorithm of Savitsky et al., (2011), that cleverly avoids having to deal with the changing dimensions of the parameter space via a joint update of the inclusion indicators of the spike-and-slab priors and the corresponding regression coefficients. We provide full details of the MCMC algorithm in the Supplementary Material. Briefly, our algorithm is composed by the following steps:
@@
    We jointly update $(\delta,\widetilde{\delta},\gamma,\beta,\widetilde{\beta})$ with a stochastic search approach. At each iteration $t=1,\dots,T$, we randomly select one $j\in\{1,\dots,p\}$ and propose to change the value of $\gamma_{j}$. If the proposed $\gamma_{j}^{*}=0$ then the only possible choice for the proposed $\beta_{j}^{*}$, $\widetilde{\beta}_{j}^{*}$, $\delta_{j}^{*}$, $\widetilde{\delta}_{j}^{*}$ is to set them all to 00. However, if we select a non-included variable, $\gamma_{j}=0$, we randomly propose to include the original variable or its knockoff. More specifically, we propose the selected coefficient using a Gaussian distribution $N(0,0.5)$ as the proposal distribution. We then accept the proposed parameters with probability $r=\min\left\{1,\rho\right\}$ with
@@
    To improve the mixing of the algorithm, we perform a within-model update step, proposing a new value for each selected coefficient using as proposal a distribution centered on the current value of $(\beta,\widetilde{\beta})$.
@@
    $\sigma^{2}\mid y,\bm{X},U,\beta,\widetilde{\beta},\bm{\Omega}\sim IG[a_{\sigma}+\frac{n}{2},b_{\sigma}+\frac{\sum_{i}(y_{i}-x_{i}(\beta+(I-\bm{\Omega}\text{diag}(s))\widetilde{\beta})-U_{i}\widetilde{\beta})^{2}}{2}]$
@@
    The update of $\bm{G}$ and $\bm{\Omega}$ is done via a block Gibbs sampler similar to the one used by Wang, (2015).
@@
    Lastly, we update the elements of the latent auxiliary variable $U$ from their full conditional distributions
    $u_{i}\mid y,\bm{X},U,\beta,\widetilde{\beta},\bm{\Omega},\sigma^{2}\sim N(\mu_{i},{\bm{\Sigma}}_{U})$
    ${\bm{\Sigma}}_{U}=\left({\bm{A}}+\frac{1}{\sigma^{2}}\widetilde{\beta}\,\widetilde{\beta}^{T}\right)^{-1}$
    $\mu_{i}={\bm{\Sigma}}_{U}\left(\frac{1}{\sigma^{2}}\left(y_{i}\widetilde{\beta}-\widetilde{\beta}\,\beta^{T}x_{i}-\widetilde{\beta}\widetilde{\beta}^{T}\Gamma_{i}\right)\right)$
    $\Gamma=X(I-{\bm{\Omega}}\text{diag}(s)),$
    for $i=1,\dots,n$.
@@ 2.3 Controlling the false discovery rate
In this section, we demonstrate that our model effectively performs variable selection while controlling the Bayesian FDR. In Bayesian frameworks, the conventional definition of the false discovery rate is not directly applicable, as both the set of active and inactive variables are treated as random variables. To address this, it is customary to rely on a notion of Bayesian FDR, as discussed for example by Whittemore, (2007), where it is defined as the expected value of the false discovery proportion conditioned on the observed data, $\bm{D}$, as
$BFDR(\mathcal{S})=\mathbb{E}\left[\frac{|\{j\in\mathcal{S}:X_{j}\mbox{$\perp\!\!\!\perp\,$}Y\}|}{|\mathcal{S}|\lor 1}\;\bigg{|}\;\bm{D}\right],$ (15)
@@
Let the prior distribution for the knockoff latent variable $U$ be such that pairwise exchangeability holds, i.e.
@@
Let the distribution of the response variable $Y$ follow the model in Equation (9) with prior distribution on the regression coefficient invariant to swaps, i.e.
@@
where $W(\beta,\widetilde{\beta})$ is an antisymmetric function, controls the $BFDR$ at level $q$.
@@
Let $r=(r_{1},\dots,r_{p})\in\{0,1\}^{p}$ be a random vector such that $r_{j}=1\iff X_{j}\mbox{$\perp\!\!\!\perp\!\!\!\!\!\!\setminus\;$}Y\mid X_{-j}$ with $X_{-j}=(X_{1},\dots,X_{j-1},X_{j+1},\dots,X_{p})$. Under a compound loss function (Muller et al.,, 2006), the Bayesian FDR can be written as
@@
To complete the proof, we need to show that an invariant prior on $(\beta,\widetilde{\beta})$ leads to a $W$ that satisfies the sign-flip property. This result can be obtained using a similar argument as in Theorem 1 of Gu and Yin, (2021). We report the proof in the Supplementary Material for completeness. ∎
@@
^$\widehat{\mathbb{P}}\left[r_{j}=0\mid\bm{D}\right]\leq 1-\frac{1}{T}\left(\sum_{t=1}^{T}\mathbbm{1}_{\{W_{j}^{(t)}>0\}}-\mathbbm{1}_{\{W_{j}^{(t)}<0\}}\right),$ (18)
where $T$ is the number of posterior samples. A simple algorithm to construct the set $\widehat{\mathcal{S}}$ required by Proposition 1 is to compute the upper bound in Equation (18) for each covariate and sort them such that the bounds are in ascending order. Then, add variables to the set until the average of the upper bounds is below $q$. The final selection is not sensitive to the choice of $W_{j}$. We adopt the typical function
$W_{j}=\Big{|}\beta_{j}\Big{|}-\left|\widetilde{\beta_{j}}\right|$
in all applications of this paper. We remark that our model performs variable selection using an upper bound on the posterior probability of non-inclusion. This strategy allows us to overcome issues with strong correlation structure among the covariates, which typically affect the state-of-art variable selection methods that rely on the estimated marginal posterior probabilities of inclusion, as empirically shown in our simulation study. Our theoretical results, indeed, show that the bound used for our variable selection procedure is not affected by the dependency structure of the covariates.
Although Proposition 1 shows that our method controls the BFDR at a threshold $q$, it assumes that the latent variables $\widetilde{X}$ are valid knockoffs a priori. As discussed by Candès et al., (2018), the construction of valid model-X knockoffs requires perfect knowledge of the covariates’ distribution. However, our proposed model assumes that the covariates $X$ follow a Gaussian distribution, with an unknown precision matrix that must be estimated from the data. Barber et al., (2020) prove that the knockoff procedure maintains FDR control at a desired threshold $q$ if the precision matrix of the Gaussian model is correctly estimated. Here, we show that, under mild regularity conditions, the priors chosen to estimate the Gaussian graphical model for the covariates yield accurate posterior estimates of the precision matrix. Thus, drawing on similar arguments as in Barber et al., (2020), we can conclude that our procedure controls both the frequentist FDR and, consequently, the expected BFDR (Guindani et al.,, 2009) in large samples, even if the precision matrix is unknown. More precisely, we are able to prove that the posterior convergence rate of our estimator is the same as the one of the Bayesian graphical lasso. To show such result, we first need to prove that our prior distribution assigns enough probability mass to values “close” to the true one.
@@
Let the true precision matrix ${\bm{\Omega}}_{0}\in\mathcal{U}(\epsilon_{0},l)=\{{\bm{\Omega}}:|\{(i,j):1\leq i<j\leq p,\omega_{ij}\neq 0\}|\leq l,0<\epsilon_{0}\leq\text{eig}({\bm{\Omega}})_{1}<\text{eig}({\bm{\Omega}})_{p}\leq\epsilon_{0}^{-1}<\infty\}$ for some $0<\epsilon_{0}<\infty$ and $20\leq l\leq p(1-p)/$ 2. Also assume that the prior distributions for the precision matrix $\Omega$ and the adjacency matrix $\bm{G}$ are:
@@
Using the prior concentration result stated in Proposition 2, we can investigate the posterior convergence rate of the graphical parameters. Proposition 3 shows how the posterior convergence rate of our model is equal to the convergence rate of the graphical lasso (Friedman et al.,, 2008) and the Bayesian graphical lasso (Park and Casella,, 2008). Note that, in Proposition 3, the prior distribution for the precision matrix in Equation (19) is slightly different from the one used in our model in Equation (11) since it includes the probability $\mathbb{P}[\bar{R}\geq|\bm{G}|]$ to penalize denser graphs. However, in the context of MCMC algorithm to sample from posterior distribution, such probability has little to no effect. The ratio of these probabilities is, indeed, close or superior to 1 for most proposal distributions.
@@
Let $X^{(n)}=(X_{1},\dots,X_{n})$ be a random sample from the p-dimensional Gaussian distribution with mean 00 and precision matrix ${\bm{\Omega}}_{0}\in\mathcal{U}(\epsilon_{0},l)=\{{\bm{\Omega}}:|\{(i,j):1\leq i<j\leq p,\omega_{ij}\neq 0\}|\leq l,0<\epsilon_{0}\leq\text{eig}({\bm{\Omega}})_{1}<\text{eig}({\bm{\Omega}})_{p}\leq\epsilon_{0}^{-1}<\infty\}$ for some $0<\epsilon_{0}<\infty$ and $20\leq l\leq p(1-p)/$ 2. Also assume that the prior distributions are:
@@
Proposition 3 shows that, with sufficiently small prior variances, the posterior mean of the precision matrix ${\bm{\Omega}}$ converges to the true one with the same rate as the convergence rate of graphical lasso (Rothman et al.,, 2008). Therefore, employing similar arguments as in Barber et al., (2020), for a sufficiently large sample size, our model controls the frequentist FDR and, consequently, the expected BFDR under a threshold $q$ without the knowledge of the true precision matrix ${\bm{\Omega}}$.

### v3

@@ 2.2.1 Extensions to non-Gaussian responses
where $T^{*}_{i}=\min(T_{i},c_{i})$ and $\kappa_{i}=\mathbbm{1}{(T_{i}\leq c_{i})}$ is the censoring indicator. Conditional on $X$, the latent response $y_{i}$ then follows the same baseline Gaussian regression model as in Equation (5). Therefore, the parameter-expanded knockoff construction described above can be applied directly to the latent response formulation of the AFT model, leading to a corresponding complete-data representation.
@@ 2.2.2 Sparsity inducing prior for covariates’ graphical model
We regard the Gaussian distribution (4) on the covariates as a graphical model and impose sparsity in the associated graph structure. Under the assumption of joint Gaussian distribution, conditional independence implies zero values in the off-diagonal elements of the precision matrix $\bm{\Omega}=\bm{\Sigma}^{-1}$. When $\bm{\Omega}$ is sparse, also the matrix $\bm{A}$ in Equation (2) is sparse, facilitating the sampling of the knockoff variables under the pairwise exchangeability property. In addition, promoting sparsity in the precision matrix $\bm{\Omega}$ greatly improves computational efficiency and improves practical interpretability across various application domains. Here, we achieve this by imposing a continuous spike-and-slab prior (Wang, 2015) that combines a mixture prior distribution on the off-diagonal elements $\omega_{jj^{\prime}}$ with an exponential distribution on the diagonal ones $\omega_{jj}$ ($j=1,\ldots,p$). Let $\mathcal{G}$ be the undirected graph representing the conditional independences among the covariates, and let $\bm{G}$ be the adjacency matrix with elements $g_{jj^{\prime}}\in\{0,1\}$, for $j,j^{\prime}=1,\ldots,p$ and $j\neq j^{\prime}$, with $1$ representing the presence of an edge between nodes $(j,j^{\prime})$. Assuming the parameters in $\bm{\Omega}$ independent a priori, the prior distribution can be written as
@@ 2.2.3 Priors for variable selection
For the regression coefficients, we adapt the spike-and-slab prior introduced in the supplementary material of Candès et al. (2018). Specifically, we modify their formulation by allowing each variable’s inclusion probability to depend on its number of neighbors in the estimated covariate graph, as detailed below. For simplicity of exposition, we formulate the prior specification and variable-selection mechanism under the augmented conditional representation in Equation (8); analogous arguments apply after marginalizing over the latent knockoff layer, yielding the observed-data model in Equation (9). We first specify the joint distribution
@@
with $j=1,\ldots,p$. Here, $h_{\beta}$ is a positive scaling hyperparameter and $\gamma=(\gamma_{1},\dots,\gamma_{p})^{T}\in\{0,1\}^{p}$ is a $p$-dimensional vector of binary indicators such that $\gamma_{j}=1$ if at least one of the two coefficients, $\beta_{j}$ and $\widetilde{\beta}_{j}$, is non-zero in the expanded representation, and $\gamma_{j}=0$ if both are zero. By introducing two additional latent indicator vectors, $\delta=(\delta_{1},\dots,\delta_{p})$ and $\widetilde{\delta}=(\widetilde{\delta}_{1},\dots,\widetilde{\delta}_{p})$, such that $\gamma=\delta+\widetilde{\delta}$, the joint prior in Equation (12) can be reformulated as a mixture of independent spike-and-slab priors on $\beta_{j}$ and $\tilde{\beta}_{j}$ as
@@
Allowing both coefficients to be non-zero simultaneously would weaken the intended one-to-one competition between them and make the resulting feature statistic harder to interpret. Therefore, we do not allow both indicators to assume value $1$ and choose a multinomial distribution of the type
$\left(\delta_{j},\widetilde{\delta}_{j}\right)\mid\gamma$ $\sim\begin{cases}(0,0)\quad\text{ w.p. }\mathbb{P}[\gamma_{j}=0]\\ (0,1)\quad\text{ w.p. }\mathbb{P}[\gamma_{j}=1]/2\\ (1,0)\quad\text{ w.p. }\mathbb{P}[\gamma_{j}=1]/2.\end{cases}$ (13)
Furthermore, we let the probability of success depend on the graphical model on the covariates, similar to Peterson et al. (2016), via an Ising prior of the type
@@
We note that Candès et al. (2018) adopt prior in Equation (12) in a fixed design Bayesian regression with augmented covariates matrix $\bm{X},\widetilde{\bm{X}}$] where $\widetilde{\bm{X}}$ is sampled once from Equation (3). On the contrary, we consider the knockoffs as latent variables and update them in the posterior sampling along with the other parameters.
@@ 2.2.4 Markov Chain Monte Carlo for posterior sampling
[…]
This distribution is not analytically tractable, and we therefore resort to a Markov Chain Monte Carlo (MCMC) algorithm (Brooks et al., 2011). In particular, we implement an efficient Metropolis within Gibbs scheme that employs a data augmentation (DA) scheme based on the latent variable $U$ introduced in equation (7) and the complete-data representation of our model detailed in Section 2.2. Specifically, since it is not possible to obtain tractable full conditional distributions under the marginal model $p(Y\mid X,\beta,\tilde{\beta},\Omega,\sigma^{2})$, we work with the augmented representation $p(Y\mid X,U,\beta,\tilde{\beta},\Omega,\sigma^{2})$ and then update $U$ from its full conditional distribution. We also incorporate the Add-Delete Metropolis-Hastings algorithm of Savitsky et al. (2011), that cleverly avoids having to deal with the changing dimensions of the parameter space via a joint update of the inclusion indicators of the spike-and-slab priors and the corresponding regression coefficients. We provide full details of the MCMC algorithm in the Supplementary Material. Briefly, our algorithm is composed by the following steps:
@@
[…]
@@
[…]
@@
    $\sigma^{2}\mid y,\bm{X},U,\beta,\widetilde{\beta},\bm{\Omega}\sim\;IG\Biggl($ $a_{\sigma}+\frac{n}{2},$
    $b_{\sigma}+\frac{1}{2}\sum_{i=1}^{n}\Bigl(y_{i}-x_{i}^{T}\bigl[\beta+\bigl(I-\bm{\Omega}\operatorname{diag}(s)\bigr)\widetilde{\beta}\bigr]-u_{i}^{T}\widetilde{\beta}\Bigr)^{2}\Biggr).$
@@
    We update $(G,\bm{\Omega})$ from $p(G,\bm{\Omega}\mid X)$ via a block Gibbs sampler similar to the one used by Wang (2015).
@@
    Lastly, we update the rows $u_{i}$ of the latent auxiliary variable $U$ from their full conditional distributions
    $u_{i}\mid y,X,\beta,\widetilde{\beta},\Omega,\sigma^{2}\sim N(\mu_{i},\Sigma_{U}),$
    where
    $\Sigma_{U}=\left(A^{-1}+\frac{1}{\sigma^{2}}\widetilde{\beta}(\widetilde{\beta})^{T}\right)^{-1},$
    and
    $\mu_{i}=\Sigma_{U}\frac{1}{\sigma^{2}}\widetilde{\beta}\left(y_{i}-x_{i}^{T}\left[\beta+\left(I-\Omega\operatorname{diag}(s)\right)\widetilde{\beta}\right]\right),\qquad i=1,\dots,n.$
@@ 2.3 Controlling the false discovery rate
In this section, we demonstrate that our model effectively performs variable selection while controlling the Bayesian FDR. In Bayesian frameworks, the conventional definition of the false discovery rate is not directly applicable, as both the set of active and inactive variables are treated as random variables. To address this, it is customary to rely on a notion of Bayesian FDR, as discussed for example by Whittemore (2007), where it is defined as the expected value of the false discovery proportion conditioned on the observed data, $\bm{D}$, as
$BFDR(\mathcal{S})=\mathbb{E}\left[\frac{|\{j\in\mathcal{S}:X_{j}\mbox{$\perp\!\!\!\perp\,$}Y{\mid X_{-j}}\}|}{|\mathcal{S}|\lor 1}\;\bigg|\;\bm{D}\right],$ (16)
@@
Let the latent knockoff matrix $\tilde{X}$ be induced from the augmentation variable $U$ through Equation (7), and assume that the resulting pair (X,X~)($\tilde{X}$) is pairwise exchangeable, i.e.
@@
[…]
@@
[…]
@@
Let $r=(r_{1},\dots,r_{p})\in\{0,1\}^{p}$ be a random vector such that $r_{j}=1\iff X_{j}\mbox{$\perp\!\!\!\perp\!\!\!\!\!\!\setminus\;$}Y\mid X_{-j}$ with $X_{-j}=(X_{1},\dots,X_{j-1},X_{j+1},\dots,X_{p})$. Under a compound loss function (Muller et al., 2006), the Bayesian FDR can be written as
@@
Since $\mathbb{P}\left[r_{j}=0\mid\bm{D}\right]\,=\,1-\mathbb{P}[r_{j}=1\mid\bm{D}]$, the inequality in Equation (17) provides an upper bound for $\mathbb{P}\left[r_{j}=0\mid\bm{D}\right]$, i.e.,
@@
$\widehat{\mathbb{P}}\left[r_{j}=0\mid\bm{D}\right]\leq 1-\frac{1}{T}\left(\sum_{t=1}^{T}\mathbbm{1}_{\{W_{j}^{(t)}>0\}}-\mathbbm{1}_{\{W_{j}^{(t)}<0\}}\right),$ (19)
where $T$ is the number of posterior samples. A simple algorithm to construct the set $\widehat{\mathcal{S}}$ required by Proposition 1 is to compute the upper bound in Equation (19) for each covariate and sort them such that the bounds are in ascending order. Then, add variables to the set until the average of the upper bounds is below $q$. Proposition 1 only requires $W_{j}$ to be antisymmetric. In practice, a choice of $W_{j}$ is needed to use the model. Many proposals are available in the literature (Barber and Candès, 2015; Spector and Janson, 2022). We adopt the simplest function
[…]
in all applications of this paper, since we found that the final selection is not sensitive to the choice of $W_{j}$.
We remark that our model performs variable selection using an upper bound on the posterior probability of non-inclusion. This strategy allows us to overcome issues with strong correlation structure among the covariates, which typically affect the state-of-art variable selection methods that rely on the estimated marginal posterior probabilities of inclusion, as empirically shown in our simulation study. Our theoretical results, indeed, show that the bound used for our variable selection procedure is not affected by the dependency structure of the covariates.
Although Proposition 1 shows that our method controls the BFDR at a threshold $q$, it assumes that the latent variables $\widetilde{X}$ are valid knockoffs a priori. As discussed by Candès et al. (2018), the construction of valid model-X knockoffs requires perfect knowledge of the covariates’ distribution. However, our proposed model assumes that the covariates $X$ follow a Gaussian distribution, with an unknown precision matrix that must be estimated from the data. Barber et al. (2020) prove that the knockoff procedure maintains FDR control at a desired threshold $q$ if the precision matrix of the Gaussian model is correctly estimated. Here, we show that, under mild regularity conditions, the priors chosen to estimate the Gaussian graphical model for the covariates yield accurate posterior estimates of the precision matrix. More precisely, we are able to prove that the posterior convergence rate of our estimator is the same as the one of the Bayesian graphical lasso. To show such result, we first need to prove that our prior distribution assigns enough probability mass to values “close” to the true one.
@@
Let the true precision matrix ${\bm{\Omega}}_{0}\in\mathcal{U}(\epsilon_{0},l)=\{{\bm{\Omega}}:|\{(i,j):1\leq i<j\leq p,\omega_{ij}\neq 0\}|\leq l,0<\epsilon_{0}\leq\text{eig}({\bm{\Omega}})_{1}<\text{eig}({\bm{\Omega}})_{p}\leq\epsilon_{0}^{-1}<\infty\}$ for some $0<\epsilon_{0}<\infty$ and $0\leq l\leq p(1-p)/2$. Also assume that the prior distributions for the precision matrix $\Omega$ and the adjacency matrix $\bm{G}$ are:
@@
Using the prior concentration result stated in Proposition 2, we can investigate the posterior convergence rate of the graphical parameters. Proposition 3 shows how the posterior convergence rate of our model is equal to the convergence rate of the graphical lasso (Friedman et al., 2008) and the Bayesian graphical lasso (Park and Casella, 2008). Note that, in Proposition 3, the prior distribution for the precision matrix in Equation (20) is slightly different from the one used in our model in Equation (11) since it includes the probability $\mathbb{P}[\bar{R}\geq|\bm{G}|]$ to penalize denser graphs. However, in the context of MCMC algorithm to sample from posterior distribution, such probability has little to no effect. The ratio of these probabilities is, indeed, close or superior to 1 for most proposal distributions.
@@
Let $X^{(n)}=(X_{1},\dots,X_{n})$ be a random sample from the p-dimensional Gaussian distribution with mean 0 and precision matrix ${\bm{\Omega}}_{0}\in\mathcal{U}(\epsilon_{0},l)=\{{\bm{\Omega}}:|\{(i,j):1\leq i<j\leq p,\omega_{ij}\neq 0\}|\leq l,0<\epsilon_{0}\leq\text{eig}({\bm{\Omega}})_{1}<\text{eig}({\bm{\Omega}})_{p}\leq\epsilon_{0}^{-1}<\infty\}$ for some $0<\epsilon_{0}<\infty$ and $0\leq l\leq p(1-p)/2$. Also assume that the prior distributions are:
@@
Proposition 3 shows that, with sufficiently small prior variances, the posterior mean of the precision matrix $\bm{\Omega}$} converges to the true one with the same rate as the convergence rate of graphical lasso (Rothman et al., 2008). Therefore, employing similar arguments as in Barber et al. (2020), we can show that our model controls the frequentist FDR up to an inflation factor that depends on how well the conditional distribution of $X$ (and hence the knockoff mechanism) is approximated. More specifically, we establish that, for a fixed nominal target level $q\in(0,1)$, the frequentist FDR is controlled at level $qe^{\eta_{n}}$, where $\eta_{n}$ depends on the posterior contraction rate of $\bm{\Omega}$} around the true precision matrix ${\bm{\Omega}}_{0}$ and the sparsity of ${\bm{\Omega}}_{0}$. In high-dimensional settings, only an inflated FDR control can be achieved under the conditions of Proposition 3, similarly as in Barber et al. (2020). Tighter control may be possible under additional structural assumptions (e.g., bounded maximum degree together with $\ell_{\infty}$ contraction of the posterior). More specifically, we can prove the following result:
Proposition 4.
Let $X_{1},\ldots,X_{n}\stackrel{{\scriptstyle iid}}{{\sim}}N_{p}\left(0,\Omega_{0}^{-1}\right)$ with $\Omega_{0}\succ 0$, and assume that $\lambda_{\min}(\Omega_{0})\geq m_{0}>0$. Assume also that $\log p=o(n)$ and that Proposition 3 holds with contraction rate $\varepsilon_{n}$, with $\varepsilon_{n}\to 0$. Let $\widehat{S}$ be as in Proposition 1, and let
$\mathbb{E}^{*}[\cdot]:=\mathbb{E}_{X^{(n)}\sim\mathbb{P}_{0}}\,\mathbb{E}_{\Omega\sim\Pi(\cdot\mid X^{(n)})}\,\mathbb{E}_{\tilde{X}\mid\Omega,X^{(n)}}[\cdot],$
denote the marginal expectation with respect to the joint law of the $X$’s, the posterior of $\Omega$, and the knockoff generation. Then there exists a sufficiently large constant $M>0$ such that
$\mathbb{E}^{*}\!\left[\mathrm{FDP}(\widehat{S})\right]\;\leq\;q\,e^{\eta_{n}}+o(1),$
where $q\in(0,1)$ and $\eta_{n}:=\Big(2M^{2}n\varepsilon_{n}^{2}+4M\varepsilon_{n}\sqrt{n\log p}\Big)(1+o(1)).$